# Anomaly induced transport from symmetry breaking in holography

### v1

@@ 1 Introduction
In recent times, the relationship between conformal field theories (CFT) and gravity theories on specific backgrounds has been extensively explored. The most prominent and thoroughly studied example of this correspondence is the one between ${\mathcal{N}}=4$ supersymmetric Yang-Mills (SYM) theory in the large $N$ limit and type IIB string theory in the AdS5 $\times$ S5 space Maldacena:1997re; Witten:1998qj. When the field theory is at strong coupling, the string theory simplifies to classical supergravity, making it possible to compute all correlation functions in the field theory.
@@
Anomaly induced transport phenomena are extremely active areas of research that arise in the context of relativistic field theories of chiral fermions, going beyond the scope of perturbation theory Kharzeev:2015znc; Bastianelli:2006rx. These anomalies have been instrumental in shaping the modern understanding of relativistic hydrodynamics Landsteiner:2012kd. Quantum anomalies, such as the chiral anomaly, give rise to novel transport phenomena such as the chiral magnetic (CME) and the chiral vortical (CVE) effect that arise at finite temperature and chemical potential. The CME is observed in systems with chiral imbalance, and it is responsible for the generation of an electric current induced by an applied magnetic field and parallel to it Fukushima:2008xe, while the CVE is the generation of an electric current induced by a vortex in the medium and parallel to the vorticity vector. These are examples of anomaly induced transport phenomena, meaning that quantum corrections to classical symmetries are the ultimate responsible of them, and they represent fascinating phenomena arising in certain quantum systems. Moreover, an important property of these phenomena is that the anomaly induced currents are non-dissipative, i.e. they don’t contribute to entropy production Son:2009tf.
@@
The anomaly-induced effects have practical implications in high-energy physics and in condensed matter systems. In high energy physics the search for signatures of anomaly induced transport in heavy ion collisions is an on-going endeavor and has culminated in the iso-bar program at RHIC and the LHC STAR:2009wot; Kharzeev:2019zgg. These collisions could lead to event-by-event parity violation, which is a significant departure from the expected symmetry properties of the system. The CVE can also be important in heavy ion collisions especially due to the proven presence of large vorticity STAR:2017ckg. In condensed matter physics the CME is at the origin of the observed large enhancement of the longitudinal conductivity of Weyl and Dirac semi-metals subject to a magnetic field Li:2014bha. The CVE has so far played a minor role in condensed matter physics.
@@
In recent years, these anomalous effects have been explored using holographic methods, providing significant insights Bhattacharyya:2007vjd; Erdmenger:2008rm; Landsteiner:2011iq; Rai:2023nxe; Gallegos:2024qxo. One such study is Megias:2016ery, where a holographic model with a pure Chern-Simons term was considered. This work examined the role of the axial anomaly in the hydrodynamics of relativistic fluids in the presence of an external electric field. In this study, the anomalous conductivities were computed using Kubo formulae. A particularly interesting result from this research is the identification of a novel phenomenon where anomalies can induce transport effects not only in anomalous currents but also in non-anomalous ones, i.e., those with a vanishing divergence, when symmetries are explicitly broken. This phenomenon was studied at strong coupling in a holographic Einstein-Maxwell model in five dimensions, supplemented with a pure gauge Chern-Simons term. Symmetry breaking was introduced through a scalar field dual to an operator $\cal O$} with $\dim{\cal O}=3$. Our goal in the present work is to extend the study for other anomalous coefficients, like the chiral vortical conductivity. In addition, we will study the role played by the mixed gauge-gravitational anomaly in the generation of the non-anomalous currents. In order to achieve these goals, one should consider the following ingredients: i) make the computation including the full backreaction of the gauge fields onto the metric tensor, and ii) add a mixed gauge-gravitational Chern-Simons terms in the action of the model. More specifically we will consider a model with three symmetries, two of them can be identified with the usual vector and axial $U(1)$ symmetries of Dirac fermions. Of these only the axial symmetry is anomalous. In addition we consider a non-anomalous symmetry [^5] To be precise we note that also the vector symmetry is non-anomalous. The vector gauge field does however appear in the Chern-Simons term and this means that it has a covariant but not a consistent anomaly. In contrast the gauge field corresponding to the additional symmetry does not appear in the Chern-Simons terms. For brevity we will refer to it as non-anomalous., and a symmetry breaking scalar field that is simultaneously charged under the axial and the non-anomalous symmetry. The boundary value $M$ of the scalar field is the symmetry breaking parameter.
@@
The structure of this manuscript is organized as follows. We summarize in section 2 the main results discussed in the literature about the constitutive relations related to anomalous transport, and motivate the introduction of a non-anomalous symmetry. In section 3 we present the holographic model under consideration, the ansatz used for the solutions of the different fields, and derive the fully backreacted numerical solution for the background. The model comprises of different background parameters such as $\mu_{v}$, $\mu_{a}$ and $\mu_{w}$, which correspond to chemical potentials for the vector, axial and non-anomalous symmetry respectively. In addition to this, there are two more background parameters: the black hole temperature $T$, and $M$ which is the symmetry breaking parameter. We fix the position of the horizon in the numerical calculation. Subsequently, in section 4, we consider the perturbations on both the metric and the gauge fields on top of the background computed in section 3.2. Next, in section 5 we present the Kubo formulae associated with these conductivities. In section 6 we present our results: first, in subsection 6.1 we start by furnishing the method used to solve numerically the equations of motion of the fluctuations, followed by subsection 6.2 where we present the comparison between our numerics and the analytic results at $M=0$ to confirm the reliability of the numerical method used. In subsection 6.3, we will present the results of the conductivities controlling the response to $\vec{B}_{v}$, $\vec{B}_{a}$, $\vec{B}_{w}$ and $\vec{\Omega}$, as a function of $\mu_{v}/(\pi T)$, $\mu_{a}/(\pi T)$ and $\mu_{w}/(\pi T)$, for different values of $M/(\pi T)$. Furthermore, we present in subsection 6.4 the results of the conductivities vs. $M/(\pi T)$ in two different cases: i) fixed non-vanishing chemical potentials, and ii) vanishing chemical potentials, i.e. $\left(\mu_{v}=\mu_{a}=\mu_{w}=0\right)$. Finally, we summarize our findings and provide an outlook for future directions in section 7. Appendix A contains the equations of motion for the background and for the fluctuations, while appendix B collects the analytical results for the conductivities in the large $M$ limit.
@@ 2 Constitutive relations and (non)anomalous transport
where $\vec{B}_{b}$ is the magnetic field corresponding to symmetry $b$. The 1-loop computation of the chiral magnetic conductivity by using the Kubo formalism leads to Kharzeev:2009pj; Landsteiner:2011cp; Chowdhury:2015pba
@@
where $d_{abc}$ is the group theoretic factor related to the axial anomaly, which typically appears in the computation of the anomalous triangle diagram corresponding to three non-abelian gauge fields coupled to a chiral fermion. The subscripts $R$, $L$ stand for the contributions of right-handed and left-handed fermions. Anomalies are responsible for a non-vanishing value of the divergence of the current, that reads in this case Kimura:1969iwz
@@ 3 Holographic action
In this work, we consider a holographic model that includes both a pure gauge and mixed gauge-gravitational Chern-Simons terms in the action, which mimics the axial anomaly and gravitational contribution to the axial anomaly respectively. The action reads Landsteiner:2013aba; Megias:2016ery
@@
and they correspond, respectively, to the Gibbons-Hawking term required to have a well-defined variational problem, and a term induced by the mixed gauge-gravitational anomaly. The values of the couplings $\kappa$ and $\lambda$ are given in terms of the anomalous coefficients of the field theory as Landsteiner:2011iq
@@ 3.1 Holographic model with symmetry breaking
A key objective of this work is to examine the possibility that the constitutive relation for $\langle J_{w}^{\mu}\rangle$ receives anomalous contributions despite the symmetry being non-anomalous. To explore this possibility, we extend our model of eq. (6) by introducing explicit symmetry breaking via a scalar field, $S_{\phi}$ Megias:2016ery, so that
@@
where, $S$ is the bulk action comprising the usual Maxwell kinetic terms and both pure and mixed Chern-Simons contributions, while the additional scalar contribution is
@@
Here $\phi$ is a scalar field with a tachyonic bulk mass $m^{2}=\Delta(\Delta-4)/\ell^{2}$, and $0\leq\Delta\leq 4$. $S_{\phi}$ produces an explicit breaking of $a$ and $w$ symmetries via the scalar field $\phi$. From the AdS/CFT dictionary, the model is the holographic dual of a Conformal Field Theory (CFT) with a deformation
${\cal L}={\cal L}^{\rm CFT}+g\,{\cal O}\,$, (12)
where $\cal O$} is an operator dual of the scalar field with $\dim{\cal O}=\Delta$, and $g$ is the source of the operator with $\dim g=4-\Delta$. The near boundary expansion of $\phi$ reads
@@
where $\phi_{4-\Delta}$ is interpreted as the source $g$, and $\phi_{\Delta}$ as the condensate $\langle{\cal O}\rangle$. In the following we will choose $\Delta=3$, so that the bulk mass is $m^{2}=-3/\ell^{2}$.
@@ 3.2 Numerical solution for the Background
where we have assumed in these expressions that the gauge fields must vanish on the horizon: $S_{t}(r_{h})=0$ (S=V,A,W)\;(S=V,A,W). It is convenient to define in the following new fields $A_{\pm}\equiv A\pm W$, so that the covariant derivative writes $D_{M}\phi=\left[\partial_{M}-i(A_{-})_{M}\right]\phi$. Consequently, the chemical potentials take the form $\mu_{\pm}\equiv\mu_{a}\pm\mu_{w}$.
@@ 4 Fluctuations
where $S_{\mu}$ refers to the background solutions computed in section 3.2 for any of the fields $V$, $A_{\pm}$ (or equivalently $V$, AA and $W$), and $s_{\mu}$ are the corresponding fluctuations $v,a_{\pm}\equiv a\pm w$ (or equivalently $v$, $a$ and $w$). Then, we will follow the general procedure of Fourier mode decomposition Amado:2011zx
@@ 5 Constitutive relations and Kubo formulae
The transport coefficients for the chiral vortical and magnetic effects, $\sigma_{V}$ and $\sigma_{B}$, have been studied in a wide variety of methods which includes Kubo formulae Amado:2011zx; Landsteiner:2012kd; Chowdhury:2015pba, fluid gravity correspondence Bhattacharyya:2007vjd; Erdmenger:2008rm; Banerjee:2008th; Megias:2013joa, diagrammatic methods Manes:2012hf and the partition function formalism Banerjee:2012iz; Megias:2014mba; Jensen:2012jy; Jensen:2012jh. In this work we use Kubo formalism to study the chiral vortical effect of anomalous conductivities. We will study the possibility that, under certain circumstances, this effect might be present in non-anomalous currents as well.
@@
The Kubo formula for the chiral magnetic conductivity for the dual field theory of the holographic model involves the retarded correlators of two currents, as it was derived in Kharzeev:2009pj. By using the compact notation introduced above, the Kubo formulae for the chiral magnetic conductivities write
@@
The Kubo formula for the chiral vortical conductivity involves the energy-momentum tensor Amado:2011zx. In the present case, the Kubo formulae write
@@ 6.1 Numerical method
To solve the coupled differential equations for the fluctuations we have used the pseudo-spectral method trefethen2000spectral; boyd2000spectral. To get the solutions we expand the fluctuations as a sum of Chebyshev polynomials in $u$ direction given by
@@
where, $(a_{\pm}^{m})_{i}$, $v_{i}^{m}$ $h_{it}^{m}$ are the coefficients of the polynomials in (33) with ($i=x,y$). These polynomials are inserted into the equation of motion for the fluctuations, and the collocation points for $u=0$ to $u=1$ are chosen in Gauss-Lobatto grid. By choosing the appropriate number of grid points corresponding to the number of coefficients one can solve the set of algebraic equations in terms of these coefficients to get the solution for the fluctuations.
@@ 6.2 Results of the conductivities for $M=0$
The case $M=0$ has been studied in the literature. The analytical results in this case are (see Ref. Landsteiner:2013aba and references therein)
@@
We present in this subsection the results for the different conductivities as a function of $\mu_{v}/(\pi T)$, $\mu_{a}/(\pi T)$ and $\mu_{w}/(\pi T)$, for several values of $M/(\pi T)$. We will present the results for the conductivities controlling the response to: i) $\vec{B}_{v}$, ii) $\vec{B}_{a}$, iii) $\vec{B}_{w}$, and iv) $\vec{\Omega}$; in this order. Unless otherwise stated, we will use as fixed values of the chemical potentials the following ones: $\mu_{v}/(\pi T)=0.1$, $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$; and we will consider three different cases: Case 1: changing $\mu_{v}/(\pi T)$, Case 2: changing $\mu_{a}/(\pi T)$, and Case 3: changing $\mu_{w}/(\pi T)$. We will show the results obtained by using the numerical method of sec. 6.1 in the range $0\leq M/(\pi T)\leq 2$. For comparison, we will also display the results obtained in the large $M$ limit by using the analytical formulas that will be derived in sec. 6.4.3 and appendix B. As we will see, all the conductivities show in general a pronounced dependence with the mass parameter $M$.
Figure 5: Case 1: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
6.3.1 Conductivities controlling the response to $\vec{B}_{v}$
  • Case 1: The results are displayed in fig. 5. We can see that $\sigma_{av}^{B}$, $\sigma_{wv}^{B}$ and $\sigma_{\varepsilon v}^{B}$ grow linearly with $\mu_{v}/(\pi T)$. With the increase $M$ the slopes of $\sigma_{av}^{B}$ and $\sigma_{\varepsilon v}^{B}$ decrease, while the slope of $\sigma_{wv}^{B}$ increases. The conductivity $\sigma_{vv}^{B}$ turns out to be independent of $\mu_{v}/(\pi T)$ independently of the value of $M$, exactly as in the absence of symmetry breaking, cf. eq. (37) Landsteiner:2013aba.
    Figure 6: Case 2: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 7: Case 3: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
6.3.2 Conductivities controlling the response to $\vec{B}_{a}$
Figure 8: Case 1: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
Figure 9: Case 2: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 10: Case 3: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
6.3.3 Conductivities controlling the response to $\vec{B}_{w}$
  • Case 2: The results are displayed in fig. 12. It can be seen that $\sigma^{B}_{vw}$ remains constant with $\mu_{a}/(\pi T)$, but its value increases steadily with $M$. The conductivities $\sigma^{B}_{aw}$ and $\sigma^{B}_{ww}$ grow linearly with $\mu_{a}/(\pi T)$, with their slopes enhanced as $M$ increases, except for the case $\sigma^{B}_{aw}$ and very large values of $M$, as it was discussed in Case 2 of sec. 6.3.2. Finally, $\sigma^{B}_{\varepsilon w}$ shows a quadratic dependence on $\mu_{a}/(\pi T)$, with an increasing curvature as $M$ increases, similar to Case 1.
  • Case 3: The results are displayed in fig. 13. The conductivity $\sigma^{B}_{vw}$ remains independent of $\mu_{w}/(\pi T)$, but it value increases steadily with $M$. The conductivities $\sigma^{B}_{aw}$ and $\sigma^{B}_{ww}$, on the other hand, grow linearly with $\mu_{w}/(\pi T)$, with their slopes enhanced as $M$ increases. The intersection observed for $\sigma^{B}_{aw}$ in the top-right panel of fig. 13 has the same discussion as in Case 3 of sec. 6.3.2. Finally, $\sigma^{B}_{\varepsilon w}$ has a linear ++ quadratic dependence on $\mu_{w}/(\pi T)$ with higher curvature for higher values of $M$. For very large values of $M$, this dependence is of the form $\sigma^{B}_{\varepsilon w}\propto(\mu_{a}+\mu_{w})^{2}$, as it can be seen in eq. (96).
Figure 11: Case 1: Plots of the conductivities $\sigma^{B}_{sw}$, where $S\in\{v,a,w,\varepsilon\}$ as a function of to $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
Figure 12: Case 2: Plots of the conductivities $\sigma^{B}_{sw}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of to $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 13: Case 3: Plots of the conductivities $\sigma^{B}_{sw}$, where $s\in\{v,a,w,\varepsilon\}$ as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
6.3.4 Conductivities controlling the response to $\vec{\Omega}$
@@
In this subsection we will show the results for the behavior of the conductivities as a function of $M/(\pi T)$ in two different cases: i) fixed non-vanishing chemical potentials, and ii) vanishing chemical potentials.

### v2

@@ 1 Introduction
In recent times, the relationship between conformal field theories (CFT) and gravity theories on specific backgrounds has been extensively explored. The most prominent and thoroughly studied example of this correspondence is the one between ${\mathcal{N}}=4$ supersymmetric Yang-Mills (SYM) theory in the large $N$ limit and type IIB string theory in the AdS5 $\times$ S5 space Maldacena (1998); Witten (1998). When the field theory is at strong coupling, the string theory simplifies to classical supergravity, making it possible to compute all correlation functions in the field theory.
@@
Anomaly induced transport phenomena are extremely active areas of research that arise in the context of relativistic field theories of chiral fermions, going beyond the scope of perturbation theory Kharzeev et al. (2016); Bastianelli and van Nieuwenhuizen (2006). These anomalies have been instrumental in shaping the modern understanding of relativistic hydrodynamics Landsteiner et al. (2013). Quantum anomalies, such as the chiral anomaly, give rise to novel transport phenomena such as the chiral magnetic (CME) and the chiral vortical (CVE) effect that arise at finite temperature and chemical potential. The CME is observed in systems with chiral imbalance, and it is responsible for the generation of an electric current induced by an applied magnetic field and parallel to it Fukushima et al. (2008), while the CVE is the generation of an electric current induced by a vortex in the medium and parallel to the vorticity vector. These are examples of anomaly induced transport phenomena, meaning that quantum corrections to classical symmetries are the ultimate responsible of them, and they represent fascinating phenomena arising in certain quantum systems. Moreover, an important property of these phenomena is that the anomaly induced currents are non-dissipative, i.e. they don’t contribute to entropy production Son and Surowka (2009).
@@
The anomaly-induced effects have practical implications in high-energy physics and in condensed matter systems. In high energy physics the search for signatures of anomaly induced transport in heavy ion collisions is an on-going endeavor and has culminated in the iso-bar program at RHIC and the LHC Abelev and others (2009); Kharzeev and Liao (2019). These collisions could lead to event-by-event parity violation, which is a significant departure from the expected symmetry properties of the system. The CVE can also be important in heavy ion collisions especially due to the proven presence of large vorticity Adamczyk and others (2017). In condensed matter physics the CME is at the origin of the observed large enhancement of the longitudinal conductivity of Weyl and Dirac semi-metals subject to a magnetic field Li et al. (2016); Gooth and others (2017). Similar anomalous chiral transport effects have been studied also in chiral superfluids with vorticity Bevan et al. (1997); Nissinen and Volovik (2022).
@@
In recent years, these anomalous effects have been explored using holographic methods, providing significant insights Bhattacharyya et al. (2008); Erdmenger et al. (2009); Landsteiner et al. (2011a); Rai and Megías (2023); Gallegos et al. (2024). One such study is Megías (2017), where a holographic model with a pure Chern-Simons term was considered. This work examined the role of the axial anomaly in the hydrodynamics of relativistic fluids in the presence of an external electric field. In this study, the anomalous conductivities were computed using Kubo formulae. A particularly interesting result from this research is the identification of a novel phenomenon where anomalies can induce transport effects not only in anomalous currents but also in non-anomalous ones, i.e., those with a vanishing divergence, when symmetries are explicitly broken. This phenomenon was studied at strong coupling in a holographic Einstein-Maxwell model in five dimensions, supplemented with a pure gauge Chern-Simons term. Symmetry breaking was introduced through a scalar field dual to an operator $\cal O$} with $\dim{\cal O}=3$. Our goal in the present work is to extend the study for other anomalous coefficients, like the chiral vortical conductivity. In addition, we will study the role played by the mixed gauge-gravitational anomaly in the generation of the non-anomalous currents. In order to achieve these goals, one should consider the following ingredients: i) make the computation including the full backreaction of the gauge fields onto the metric tensor, and ii) add a mixed gauge-gravitational Chern-Simons terms in the action of the model. More specifically we will consider a model with three symmetries, two of them can be identified with the usual vector and axial $U(1)$ symmetries of Dirac fermions. Of these only the axial symmetry is anomalous. In addition we consider a non-anomalous symmetry [^5] To be precise we note that also the vector symmetry is non-anomalous. The vector gauge field does however appear in the Chern-Simons term and this means that it has a covariant but not a consistent anomaly. In contrast the gauge field corresponding to the additional symmetry does not appear in the Chern-Simons terms. For brevity we will refer to it as non-anomalous., and a symmetry breaking scalar field that is simultaneously charged under the axial and the non-anomalous symmetry. The boundary value $M$ of the scalar field is the symmetry breaking parameter.
@@
The structure of this manuscript is organized as follows. We summarize in section 2 the main results discussed in the literature about the constitutive relations related to anomalous transport, and motivate the introduction of a non-anomalous symmetry. In section 3 we present the holographic model under consideration, the ansatz used for the solutions of the different fields, and derive the fully backreacted numerical solution for the background. The model comprises of different background parameters such as $\mu_{v}$, $\mu_{a}$ and $\mu_{w}$, which correspond to chemical potentials for the vector, axial and non-anomalous symmetry respectively. In addition to this, there are two more background parameters: the black hole temperature $T$, and $M$ which is the symmetry breaking parameter. We fix the position of the horizon in the numerical calculation. Subsequently, in section 4, we consider the perturbations on both the metric and the gauge fields on top of the background computed in section 3.2. Next, in section 5 we present the Kubo formulae associated with these conductivities. In section 6 we present our results: first, in subsection 6.1 we start by furnishing the method used to solve numerically the equations of motion of the fluctuations, followed by subsection 6.2 where we present the comparison between our numerics and the analytic results at $M=0$ to confirm the reliability of the numerical method used. In subsection 6.3, we will present the results of the $\mu_{v}$-dependence of the conductivities that induce the non-anomalous current $\vec{J}_{w}$, for different values of $M/(\pi T)$. Furthermore, we present in subsection 6.4 the results of the conductivities vs. $M/(\pi T)$ in two different cases: i) fixed non-vanishing chemical potentials, and ii) vanishing chemical potentials, i.e. $\left(\mu_{v}=\mu_{a}=\mu_{w}=0\right)$. We summarize our findings and provide an outlook for future directions in section 7. Finally, appendix A contains the equations of motion for the background and for the fluctuations, appendix B collects the numerical results for all the conductivities as a function of the chemical potentials, and appendix C presents the analytical results for the conductivities in the large $M$ limit.
@@ 2 Constitutive relations and (non)anomalous transport
where $\vec{B}_{b}$ is the magnetic field corresponding to symmetry $b$. The 1-loop computation of the chiral magnetic conductivity by using the Kubo formalism leads to Kharzeev and Warringa (2009); Landsteiner et al. (2011b); Chowdhury and David (2015)
@@
where $d_{abc}$ is the group theoretic factor related to the axial anomaly, which typically appears in the computation of the anomalous triangle diagram corresponding to three non-abelian gauge fields coupled to a chiral fermion. The subscripts $R$, $L$ stand for the contributions of right-handed and left-handed fermions. Anomalies are responsible for a non-vanishing value of the divergence of the current, that reads in this case Kimura (1969)
@@ 3 Holographic action
In this work, we consider a holographic model that includes both a pure gauge and mixed gauge-gravitational Chern-Simons terms in the action, which mimics the axial anomaly and gravitational contribution to the axial anomaly respectively. The action reads Landsteiner et al. (2014); Megías (2017)
@@
and they correspond, respectively, to the Gibbons-Hawking term required to have a well-defined variational problem, and a term induced by the mixed gauge-gravitational anomaly. The values of the couplings $\kappa$ and $\lambda$ are given in terms of the anomalous coefficients of the field theory as Landsteiner et al. (2011a)
@@ 3.1 Holographic model with symmetry breaking
A key objective of this work is to examine the possibility that the constitutive relation for $\langle J_{w}^{\mu}\rangle$ receives anomalous contributions despite the symmetry being non-anomalous. To explore this possibility, we extend our model of eq. (6) by introducing explicit symmetry breaking via a scalar field, $S_{\phi}$ Megías (2017), so that
@@
where, $\mathcal{S}$ is the bulk action comprising the usual Maxwell kinetic terms and both pure and mixed Chern-Simons contributions, while the additional scalar contribution is
@@
Here $\phi$ is a scalar field with a tachyonic bulk mass $m^{2}=\Delta(\Delta-4)/\ell^{2}$, and $0\leq\Delta\leq 4$. $\mathcal{S}_{\phi}$ produces an explicit breaking of $a$ and $w$ symmetries via the scalar field $\phi$. From the AdS/CFT dictionary, the model is the holographic dual of a Conformal Field Theory (CFT) with a deformation
${\cal L}={\cal L}^{\rm CFT}+M\,{\cal O}\,$, (12)
where $\cal O$} is an operator dual of the scalar field with $\dim{\cal O}=\Delta$, and $M$ is the source of the operator with $\dim M=4-\Delta$. The near boundary expansion of $\phi$ reads
@@
where $\phi_{4-\Delta}$ is interpreted as the source $M$, and $\phi_{\Delta}$ as the condensate $\langle{\cal O}\rangle$. In the following we will choose $\Delta=3$, so that the bulk mass is $m^{2}=-3/\ell^{2}$.
@@ 3.2 Numerical solution for the Background
where we have assumed in these expressions that the gauge fields must vanish on the horizon: $S_{t}(r_{h})=0$ (S=V,A,W)\;(S=V,A,W). It is convenient to define the following new fields $A_{\pm}\equiv A\pm W$, so that the covariant derivative writes $D_{M}\phi=\left[\partial_{M}-i(A_{-})_{M}\right]\phi$. Consequently, the chemical potentials take the form $\mu_{\pm}\equiv\mu_{a}\pm\mu_{w}$.
@@ 4 Fluctuations
where $S_{\mu}$ refers to the background solutions computed in section 3.2 for any of the fields $V$, $A_{\pm}$ (or equivalently $V$, AA and $W$), and $s_{\mu}$ are the corresponding fluctuations $v,a_{\pm}\equiv a\pm w$ (or equivalently $v$, $a$ and $w$). Then, we will follow the general procedure of Fourier mode decomposition Amado et al. (2011)
@@ 5 Constitutive relations and Kubo formulae
The transport coefficients for the chiral vortical and magnetic effects, $\sigma_{V}$ and $\sigma_{B}$, have been studied in a wide variety of methods which includes Kubo formulae Amado et al. (2011); Landsteiner et al. (2013); Chowdhury and David (2015), fluid gravity correspondence Bhattacharyya et al. (2008); Erdmenger et al. (2009); Banerjee et al. (2011); Megías and Pena-Benitez (2013), diagrammatic methods Mañes and Valle (2013) and the partition function formalism Banerjee et al. (2012); Megías and Valle (2014); Jensen (2012); Jensen et al. (2012). In this work we use Kubo formalism to study the chiral vortical effect of anomalous conductivities. We will study the possibility that, under certain circumstances, this effect might be present in non-anomalous currents as well.
@@
The Kubo formula for the chiral magnetic conductivity for the dual field theory of the holographic model involves the retarded correlators of two currents, as it was derived in Kharzeev and Warringa (2009). By using the compact notation introduced above, the Kubo formulae for the chiral magnetic conductivities write
@@
The Kubo formula for the chiral vortical conductivity involves the energy-momentum tensor Amado et al. (2011). In the present case, the Kubo formulae write
@@ 6.1 Numerical method
To solve the coupled differential equations for the fluctuations we have used the pseudo-spectral method Trefethen (2000); Boyd (2001). To get the solutions we expand the fluctuations as a sum of Chebyshev polynomials in $u$ direction given by
@@
where $(a_{\pm}^{m})_{i}$, $v_{i}^{m}$ and $h_{it}^{m}$ are the coefficients of the polynomials in (33) with ($i=x,y$). These polynomials are inserted into the equation of motion for the fluctuations, and the collocation points for $u=0$ to $u=1$ are chosen in Gauss-Lobatto grid. By choosing the appropriate number of grid points corresponding to the number of coefficients one can solve the set of algebraic equations in terms of these coefficients to get the solution for the fluctuations.
@@ 6.2 Results of the conductivities for $M=0$
The case $M=0$ has been studied in the literature. The analytical results in this case are (see Ref. Landsteiner et al. (2014) and references therein)
@@
We present in this subsection the results for the conductivities as a function of the chemical potentials, for several values of $M/(\pi T)$. Here we will present only the $\mu_{v}$-dependence of the conductivities that induce the non-anomalous current $\vec{J}_{w}$, i.e. the conductivities appearing in
$\vec{J}_{w}=\sigma^{B}_{wv}\,\vec{B}_{v}+\sigma^{B}_{wa}\,\vec{B}_{a}+\sigma^{B}_{ww}\,\vec{B}_{w}+\sigma^{V}_{w}\,\vec{\Omega}\,.$ (46)
We focus momentarily in this subsection on these conductivities because they allow us to provide an explicit answer to the question posed at the end of section 2. The complete list of results for these and other conductivities is given in appendix B.
We display in fig. 5 the results of the conductivities in eq. (46) as a function of $\mu_{v}/(\pi T)$ in the range $0\leq M/(\pi T)\leq 2$. These results have been obtained by using the numerical method of sec. 6.1. For comparison, we display also the results obtained in the large $M$ limit by using the analytical formulas that will be derived in sec. 6.4.3 and appendix C. We observe that the conductivity $\sigma^{B}_{wv}$ exhibits a linear growth with $\mu_{v}/(\pi T)$, with a slope that increases as $M$ increases. On the other hand, the conductivities $\sigma^{B}_{wa}$ and $\sigma^{B}_{ww}$ remain constant with respect to $\mu_{v}/(\pi T)$, though their magnitudes shift upward with increasing $M$. Finally, we find for $\sigma^{V}_{w}$ a quadratic dependence on $\mu_{v}/(\pi T)$, with an increasing curvature as $M$ increases. Notice that all the conductivities show in general a pronounced dependence with the mass parameter $M$.
Figure 5: Plots of the conductivities $\sigma^{B}_{wv}$, $\sigma^{B}_{wa}$, $\sigma^{B}_{ww}$ and $\sigma^{V}_{w}$ as a function of to $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
@@
In this subsection we will show the results for the behaviour of the conductivities as a function of $M/(\pi T)$ in two different cases: i) fixed non-vanishing chemical potentials, and ii) vanishing chemical potentials.